Current shares outstanding $7 million
No debt
Sticks sells for $65 per share
Book value per share $20
Net income $11.5 million
Wish to expand facilities and amount of cost $30 million
And it will increase net income by $675000
The per value of the stock is $1

Under constant price-earnings ratio what will be the effect be of issuing new equity to finance the investment?
1
Expert's answer
2019-12-27T09:02:39-0500

The company has to issue 30,000,000/65 = 461,538 of new shares.

New Book Value per Share = (Pre-rights offering share value + Share Value of rights offerings)/Ex-rights outstanding shares = (7,000,000×20 + 461,538×65)/(7,000,000 + 461,538) = 22.78.

New EPS is: EPS = Income/Shares outstanding = (11,500,000 + 675,000)/(7,000,000 + 461,538) = $1.63.

Price Earnings ratio is P/E = Market Price of the share / EPS = $65/$1.63 = 39.88 times.

New share price is P = P/E×New EPS = 39.88×1.63 = $65.